OpenStudy (anonymous):

the length of a rectangle is 6 time its width. The perimeter is 98 ft. What is its length and width? the length of a rectangle is 6 time its width. The perimeter is 98 ft. What is its length and width? @Mathematics

OpenStudy (anonymous):

p=2(l+w) p=98 l=6w l=6x, w =x 98=2(6x+x) 98 = 2(7x) 98 = 14x x=7 l = 42, w = 7

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Thats alrighty.